You're are taking a multiple choice test that has five questions. Each of the questions has 3 answer choices, with 1 correct answer per question. If you guess for each one and leave nothing blank, in how many ways can you answer the questions? (and..) What is the probability that all of your answers will be INcorrect?
Answer by stanbon(75887)

You're are taking a multiple choice test that has five questions. Each of the questions has 3 answer choices, with 1 correct answer per question. If you guess for each one and leave nothing blank,
a) in how many ways can you answer the questions?
Ans: 3^5 = 243 ways
----------------------------
b) What is the probability that all of your answers will be INcorrect?
P(incorrect) = 2/3
---
P(all 5 are incorrect) = (2/3)^5 = 32/243
=================
Cheers,
Stan H.